OpenStudy (anonymous):

Simplify 1/(1+(x/y))

OpenStudy (anonymous):

Is that the answer?

zepdrix (zepdrix):

\[\large \frac{1}{1+\dfrac{x}{y}}\color{royalblue}{\left(\dfrac{y}{y}\right)}\]Multiply the fraction by y/y.\[\large \frac{y}{y+x}\]I don't believe it can be simplified any further.

OpenStudy (anonymous):

ok thanks